OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b (-5,2); 8x=9y+4

OpenStudy (anonymous):

how would you begin?

hero (hero):

I'd put the equation in terms of y first to get the slope

OpenStudy (anonymous):

good, so what did you get?

hero (hero):

That's what I want to know...What did she get....

OpenStudy (anonymous):

8x =9y +4, 9y = 8x-4, y = 8/9x - 4/9 or..... since we see here that the slop is 8/9, parallel lines have the same slope,.. and using the equation y - y=m(x-x), we get y-2=8/9(x--5), y-2=8/9x+40/9, y=8/9x+58/9... so in y=mx +b form that is the answer

OpenStudy (anonymous):

y=8/9x+58/9?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you